A 480 MHz CHIRPED-PULSE FOURIER-TRANSFORM MICROWAVE SPECTROMETER: CONSTRUC-TION AND MEASUREMENT OF THE ROTATIONAL SPECTRA OF DIVINYL SILANE AND 3,3-DIFLUOROPENTANE

\underline{DANIEL A. OBENCHAIN}, AMANDA L. STEBER, ASHLEY A. ELLIOTT, REBECCA A. PEE-BLES and SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, IL 61920; CHARLES J. WURREY, Department of Chemistry, University of Missouri-Kansas City, Kansas City, MO 64110; GAMIL A. GUIRGIS, Department of Chemistry and Biochemistry, College of Charleston, Charleston, SC 29424.

A chirped-pulse Fourier-transform microwave (CP-FTMW) spectrometer based on the original Pate design^{a} has been constructed to allow analysis of any 480 MHz region in the 7 – 18 GHz range. A 1 μs chirped-pulse (0 – 240 MHz) from an arbitrary function generator is mixed with output from a microwave synthesizer and used to polarize a supersonic gas expansion; the resulting free induction decay is collected over 20 μs and Fourier-transformed on a 500 MHz oscilloscope to produce a rotational spectrum. A variety of molecules have now been studied with this instrument and results will be presented for numerous conformers of divinyl silane (predicted μ_{total} = 0.6−0.7 D) and the more polar 3,3-difluoropentane (predicted μ_{total} = 2.5 − 2.8 D).

Two of the three possible conformers of divinyl silane were assigned (both having a C_{1}=C_{2}–Si–C_{3} dihedral angle of −120^{∘} and a C_{2}–Si–C_{3}=C_{4} dihedral of either 0^{∘} (C_{1} symmetry) or −120^{∘} (C_{2} symmetry)). For 3,3-difluoropentane, three of the four possible conformers were identified: anti-gauche (C_{1}), gauche-gauche (C_{2}) and anti-anti (C_{2v}). While rotational spectra for only the silicon isotopologues were observed for divinyl silane, measurement of the ^{13}C spectra of 3,3-difluoropentane allowed heavy atom structure determinations for the anti-gauche and gauche-gauche conformers. Initial assignments of all spectra were made on the CP-FTMW spectrometer, and a Balle-Flygare FTMW spectrometer was used to compare frequencies of measured transitions and also to provide Stark effect data. Substitution (r_{s}) and inertial fit (r_{0}) structures will be compared with computational data and instrumental details will be presented.

aG.G. Brown, B.C. Dian, K.O. Douglass, S.M. Geyer, S.T. Shipman, B.H. Pate, Rev. Sci. Instrum., 79, (2008), 053103.